Comment

Comment by turc1656

This is some great info for the less technically knowledgeable about Tor (like me!). However, I think your math in #3 is wrong.

Assuming random assignment/selection of the guards, each time one is chosen it has a 98% chance of not being "caught" by choosing an adversary's guard. Going with 50 services as you said would be .98^50=.364, meaning the chance of getting caught is 1-.364=.635 - 63.5%. This is vastly different than being nearly 100%.

parent turc1656 6 years ago

Discussion

Live from the official Hacker News API

A modern reader for Hacker News · SvelteKit + Skeleton